# Random Sperner lemma and random Brouwer fixed point theorem

### v3

@@ 2.2 A general $L^{0}$-simplicial subdivision and random Sperner lemma
[…]
Method 1. First, let us define a real-valued function $r:\Omega\rightarrow[0,+\infty)$ by $r(\omega)=\sup\{|x|:x\in U(\omega)\}$ for any $\omega\in\Omega$, then $r$ is well-defined since each $U(\omega)$ is bounded. Now, let $G_{U}^{0}=\{x^{0}:x^{0}$ is an $\mathbb{R}^{d}$-valued $\mathcal{F}$-measurable function on $\Omega$ such that $x^{0}(\omega)\in U(\omega)~{}a.s.\}$, then $|x^{0}(\omega)|\leq r(\omega)$ a.s. for each $x^{0}\in G_{U}^{0}$. Further, let $\xi^{0}=ess.sup\{|x^{0}|:x^{0}\in G_{U}^{0}\}$, see, for example, [23, 26] for the notion of an essential supremum for a set of extended real-valued random variables, then it still holds that $\xi^{0}(\omega)\leq r(\omega)$ a.s., so we can assume that $\xi^{0}$ is a real-valued random variable. Then $\bigvee\{\|x\|:x\in G_{U}\}\leq\xi$, where $\xi$ is the equivalence class of $\xi^{0}$, namely $G_{U}$ is a.s. bounded.
Method 2. Let $\hat{\mathcal{F}}$ be the completion of $\mathcal{F}$ with respect to $P$, then $Gr(U)\in\hat{\mathcal{F}}\otimes Bor(\mathbb{R}^{d})$ and hence by Theorem 4.2 of [51] there exists a sequence $\{x_{n}^{0},n\in\mathbb{N}\}$ of $\hat{\mathcal{F}}$-measurable selections of $U$ such that $U(\omega)=\overline{\{x_{n}^{0}(\omega):n\in\mathbb{N}\}}$ for each $\omega\in\Omega$. Let $r^{0}:\Omega\rightarrow[0,+\infty)$ be the function defined by $r^{0}(\omega)=\sup\{|x_{n}^{0}(\omega)|:n\in\mathbb{N}\}$ for any $\omega\in\Omega$, then $r^{0}$ is $\hat{\mathcal{F}}$-measurable and hence there exists an $\mathcal{F}$-measurable nonnegative random variable $r_{1}^{0}$ such that $r^{0}(\omega)=r_{1}^{0}(\omega)$ a.s. Further, let $r_{1}$ be the equivalence class of $r_{1}^{0}$, then it is clear that $\bigvee\{\|x\|:x\in G_{U}\}\leq r_{1}$, $G_{U}$ is again a.s. bounded. ∎
@@
For the reverse implication, for any $\xi\in L_{++}^{0}(\mathcal{F})$, let $T_{\xi}$ be the isomorphism of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ to itself, defined by $T_{\xi}(x)=\xi x$ for any $x\in L^{0}(\mathcal{F},\mathbb{R}^{d})$. Then, it is also clear that $T_{\xi}^{-1}=T_{\xi^{-1}}$. Now, let $G$ and $f$ be the same as assumed in Proposition 2.3, then there exists some $\xi\in L_{++}^{0}(\mathcal{F})$ such that $\xi^{-1}G\subseteq L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$, where $\bar{B}_{1}(\mathbb{R}^{d})=\{y\in\mathbb{R}^{d}:|y|\leq 1\}$ and ||$\cdot$|| ⋅ | stands for the Euclidean norm on $\mathbb{R}^{d}$. For brevity, we denote $L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ by $V$, then $V$ is an a.s. bounded, $\mathcal{T}_{\varepsilon,\lambda}$-closed and $L^{0}$-convex subset of $L^{0}(\mathcal{F},\mathbb{R}^{d})$. Further, let $P_{V,\xi^{-1}G}$ be the restriction of $P_{\xi^{-1}G}$ to $V$, where $P_{\xi^{-1}G}$ is the random projection of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ onto $\xi^{-1}G$, then $h:=T_{\xi^{-1}}\circ f\circ T_{\xi}\circ P_{V,\xi^{-1}G}$ is a $\sigma$-stable and $\mathcal{T}_{\varepsilon,\lambda}$-continuous mapping from $V$ to $V$, and hence $h$ is also a local and continuous-in-probability operator. Now, define a set-valued function $U:\Omega\rightarrow 2^{\mathbb{R}^{d}}$ by $U(\omega)=\bar{B}_{1}(\mathbb{R}^{d})$ for each $\omega\in\Omega$, then $U$ not only satisfies the hypothesis of Proposition 2.4 but also is weakly measurable in the sense of [51], and in particular $G_{U}$ is exactly $V$. Thus, there exists $v\in V$ such that $h(v)=v$ by Proposition 2.4, then $\xi v=f(\xi P_{V,\xi^{-1}G}(v))\in G$. Again, let $g\in G$ be such that $g=\xi v$, then $v=\xi^{-1}g\in\xi^{-1}G$, then $P_{V,\xi^{-1}G}(v)=\xi^{-1}g$, and hence $f(\xi P_{V,\xi^{-1}G}(v))=f(g)=\xi v=g$, namely $g$ is a fixed point of $f$ in $G$. ∎
@@
Let $V=L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ and $h:V\rightarrow V$ be the same as obtained in the process of the proof of Proposition 2.4 implying Proposition 2.3. If $(\Omega,\mathcal{F},P)$ is a complete probability space, then by Lemma 5.2 there exists a Carathéodory function $f:\Omega\times\bar{B}_{1}(\mathbb{R}^{d})\rightarrow\bar{B}_{1}(\mathbb{R}^{d})$ such that $h(v)=h_{f}(v)$ for any $v\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$. This observation leads us to the following interesting (or even somewhat surprising) result.
@@
    Every Carathéodory function function $f:(\Omega,\mathcal{F},P)\times\bar{B}_{1}(\mathbb{R}^{d})\rightarrow\bar{B}_{1}(\mathbb{R}^{d})$ admits a random fixed point, namely there exists a random vector $v:(\Omega,\mathcal{F},P)\rightarrow\bar{B}_{1}(\mathbb{R}^{d})$ such that $f(\omega,v(\omega))=v(\omega)$ a.s..
@@
$(1)\Rightarrow(2)$. Let $h=h_{f}$, then there exists some $x\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ such that $h_{f}(x)=x$, and hence taking $v$ to be an arbitrarily chosen representative of $x$ yields $f(\omega,v(\omega))=v(\omega)$ a.s., namely $v$ satisfies (2).
$(2)\Rightarrow(1)$. Let $\hat{\mathcal{F}}$ be the completion of $\mathcal{F}$ with respect to $P$ and further notice the fact that $L^{0}(\hat{\mathcal{F}},\bar{B}_{1}(\mathbb{R}^{d}))=L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ although $\hat{\mathcal{F}}$ and $\mathcal{F}$ may differ, then by Lemma 5.2 there exists some Carathéodory function $f:(\Omega,\hat{\mathcal{F}},P)\times\bar{B}_{1}(\mathbb{R}^{d})\rightarrow\bar{B}_{1}(\mathbb{R}^{d})$ such that $h=h_{f}$. By (1) there exists some $\hat{\mathcal{F}}$-measurable $v:\Omega\rightarrow\bar{B}_{1}(\mathbb{R}^{d})$ such that $f(\omega,v(\omega))=v(\omega)$ a.s.. Let $x$ be the equivalence class of $v$, then $h(x)=h_{f}(x)=x$, since $L^{0}(\hat{\mathcal{F}},\bar{B}_{1}(\mathbb{R}^{d}))=L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$, $x$ is also regarded as a fixed point of $h$ in $L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$. ∎
@@
Let $G$ be a $\mathcal{T}_{\varepsilon,\lambda}$-closed $L^{0}$-convex subset of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ and $T:G\rightarrow L^{0}(\mathcal{F},\mathbb{R}^{d})$ be a $\sigma$-stable $\mathcal{T}_{\varepsilon,\lambda}$-continuous mapping. Then $T$ must be a.s. sequentially continuous.
@@
Since elements of both $G$ and $L^{0}(\mathcal{F},\mathbb{R}^{d})$ are equivalence classes, we can, without loss of generality, assume that $(\Omega,\mathcal{F},P)$ is complete. Further, let $P_{G}$ be the random projection of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ onto $G$, then $T\circ P_{G}:L^{0}(\mathcal{F},\mathbb{R}^{d})\rightarrow L^{0}(\mathcal{F},\mathbb{R}^{d})$ is still $\sigma$-stable and $\mathcal{T}_{\varepsilon,\lambda}$-continuous, and hence by Lemma 5.2 there exists some Carathéodory function $f:(\Omega,\mathcal{F},P)\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ such that $T\circ P_{G}=h_{f}$, in particular $T(g)=h_{f}(g)$ for any $g\in G$, which has implied that $T$ is a.s. sequentially continuous. ∎
@@
6 The equivalence of random Brouwer fixed point theorem and random Borsuk theorem
Proposition 6.1 below shows that Theorem 2.1 is equivalent to a special case of it, namely random Brouwer fixed point theorem for random closed balls, which will provide much convenience for our research into random Borsuk theorem.
@@ 6 The equivalence of random Brouwer fixed point theorem and random Borsuk theorem
    Every $\mathcal{T}_{c}$-continuous $\sigma$-stable mapping from a $\mathcal{T}_{c}$-closed, $\sigma$-stable, a.s. bounded and $L^{0}$-convex subset of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ to itself has a fixed point.
@@
    Every $\mathcal{T}_{c}$-continuous $\sigma$-stable mapping from a random closed ball $\bar{B}(\theta,r)$ of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ to itself has a fixed point, where $r\in L^{0}_{++}(\mathcal{F})$ and $\bar{B}(\theta,r)=\{x\in L^{0}(\mathcal{F},\mathbb{R}^{d}):\|x\|\leq r\}$.
@@
    Every $\mathcal{T}_{c}$-continuous $\sigma$-stable mapping from the random closed unit ball $\bar{B}(\theta,1)$ of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ to itself has a fixed point.
@@
$(2)\Rightarrow(1)$. Let $G$ be a $\mathcal{T}_{c}$-closed, $\sigma$-stable, a.s. bounded and $L^{0}$-convex subset of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ and $f:G\rightarrow G$ be a $\mathcal{T}_{c}$-continuous $\sigma$-stable mapping. Since $G$ is a.s. bounded, we can choose an enough large $r\in L^{0}_{++}(\mathcal{F})$ such that $G\subseteq\bar{B}(\theta,r)$. Further, since $G$ is $\sigma$-stable and $\mathcal{T}_{c}$-closed, it follows immediately from [23] that $G$ is also $\mathcal{T}_{\varepsilon,\lambda}$-closed, then by Lemma 4.11 the random projection $P_{G}$ of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ onto $G$ is well defined, $\sigma$-stable and $\mathcal{T}_{c}$-continuous. Now, let $P_{\bar{B}(\theta,r),G}$ be the restriction of $P_{G}$ to $\bar{B}(\theta,r)$, then it is also $\sigma$-stable and $\mathcal{T}_{c}$-continuous since $\bar{B}(\theta,r)$ is $\sigma$-stable. Thus, it is clear that $f\circ P_{\bar{B}(\theta,r),G}$ is a $\mathcal{T}_{c}$-continuous and $\sigma$-stable mapping from $\bar{B}(\theta,r)$ to $\bar{B}(\theta,r)$, then by (2) $f\circ P_{\bar{B}(\theta,r),G}$ has a fixed point $x$. Eventually, $f(x)=x$ since $x\in G$.
$(2)\Leftrightarrow(3)$ is clear since, $T_{r}:\bar{B}(\theta,1)\rightarrow\bar{B}(\theta,r)$ defined by $T_{r}(x)=rx$ for each $x\in\bar{B}(\theta,1)$, is a $\sigma$-stable $\mathcal{T}_{c}$-homeomorphism. ∎
@@
Let $E$ and $F$ be the same as in Definition 6.2. If $h:F\rightarrow E$ is a $\sigma$-stable and bijective mapping, then it is easy to check that $h^{-1}:E\rightarrow F$ is also $\sigma$-stable. It is also clear that a $\sigma$-stable $\mathcal{T}_{c}$-homeomorphism (resp., $\mathcal{T}_{\varepsilon,\lambda}$-homeomorphism) of $F$ onto $E$ must be a $\mathcal{T}_{c}$-$r$-mapping (resp., a $\mathcal{T}_{\varepsilon,\lambda}$-$r$-mapping) of $F$ onto $E$. It follows immediately from Lemma 4.3 of [26] that a $\mathcal{T}_{\varepsilon,\lambda}$-$r$-mapping (resp., a $\mathcal{T}_{\varepsilon,\lambda}$-retraction) of $F$ onto $E$ must be a $\mathcal{T}_{c}$-$r$-mapping (resp., a $\mathcal{T}_{c}$-retraction) of $F$ onto $E$.
Proposition 6.3 below is a random version of Theorem 6.9 of [1], which is essentially a reformulation of Theorem 2.1 in a more general form.
@@
Let $g:E\rightarrow F$ and $h:F\rightarrow E$ be the same as in Definition 6.2, then $g\circ f\circ h:F\rightarrow F$ has a fixed point $z$ by Theorem 2.1. Let $x=h(z)$, then
@@
Theorem 6.4 below can be aptly called random Borsuk theorem, which is a random version of the classical Borsuk theorem [2] (see also [1, Theorem 6.12]). Here, we would like to remind the reader of the fact that $L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ in the sequel of this section is the same thing as the random closed unit ball $\bar{B}(\theta,1):=\{x\in L^{0}(\mathcal{F},\mathbb{R}^{d}):\|x\|\leq 1\}$.
@@
$L^{0}(\mathcal{F},\partial\bar{B}_{1}(\mathbb{R}^{d}))$ is not a $\mathcal{T}_{c}$-$r$-image of $L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$, where, as usual, $\bar{B}_{1}(\mathbb{R}^{d})=\{y\in\mathbb{R}^{d}:|y|\leq 1\}$, $\partial\bar{B}_{1}(\mathbb{R}^{d})=\{y\in\mathbb{R}^{d}:|y|=1\}$ and ||$\cdot$|| ⋅ | denotes the Euclidean norm on $\mathbb{R}^{d}$. In particular, $L^{0}(\mathcal{F},\partial\bar{B}_{1}(\mathbb{R}^{d}))$ is not a $\mathcal{T}_{c}$-retraction of $L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$.

### v5

@@ 2.2 A general $L^{0}$-simplicial subdivision and random Sperner lemma
For clarity, we use $|\cdot|$ for the usual Euclidean norm on $\mathbb{R}^{d}$ and reserve $\|\cdot\|$ for the $L^{0}$-norm on $L^{0}(\mathcal{F},\mathbb{R}^{d})$. We employ two methods to prove this lemma as follows. We only need to prove that $G_{U}$ is a.s. bounded since the others are obvious.
Method 1. First, define a real-valued function $r:\Omega\rightarrow[0,+\infty)$ by $r(\omega)=\sup\{|x|:x\in U(\omega)\}$ for any $\omega\in\Omega$. Then $r$ is well-defined since each $U(\omega)$ is bounded. Now, let $G_{U}^{0}=\{x^{0}:x^{0}$ is an $\mathbb{R}^{d}$-valued $\mathcal{F}$-measurable function on $\Omega$ such that $x^{0}(\omega)\in U(\omega)~a.s.\}$. Then $|x^{0}(\omega)|\leq r(\omega)$ a.s. for each $x^{0}\in G_{U}^{0}$. Further, let $\xi^{0}=ess.sup\{|x^{0}|:x^{0}\in G_{U}^{0}\}$, see, for example, [24, 27] for the notion of an essential supremum for a set of extended real-valued random variables. Then it still holds that $\xi^{0}(\omega)\leq r(\omega)$ a.s., so we can assume that $\xi^{0}$ is a real-valued random variable. Then $\bigvee\{\|x\|:x\in G_{U}\}\leq\xi$, where $\xi$ is the equivalence class of $\xi^{0}$, namely, $G_{U}$ is a.s. bounded.
Method 2. Let $\hat{\mathcal{F}}$ be the completion of $\mathcal{F}$ with respect to $P$. Then the graph $Gr(U)\in\hat{\mathcal{F}}\otimes Bor(\mathbb{R}^{d})$ and hence by [55, Theorem 4.2] there exists a sequence $\{x_{n}^{0},n\in\mathbb{N}\}$ of $^\hat{\mathcal{F}}$-measurable selections of $U$ such that $U(\omega)=\overline{\{x_{n}^{0}(\omega):n\in\mathbb{N}\}}$ for each $\omega\in\Omega$. Let $r^{0}:\Omega\rightarrow[0,+\infty)$ be the function defined by $r^{0}(\omega)=\sup\{|x_{n}^{0}(\omega)|:n\in\mathbb{N}\}$ for any $\omega\in\Omega$, then $r^{0}$ is $^\hat{\mathcal{F}}$-measurable and hence there exists an $\mathcal{F}$-measurable nonnegative random variable $r_{1}^{0}$ such that $r^{0}(\omega)=r_{1}^{0}(\omega)$ a.s. Further, let $r_{1}$ be the equivalence class of $r_{1}^{0}$. Then it is clear that $\bigvee\{\|x\|:x\in G_{U}\}\leq r_{1}$; $G_{U}$ is again a.s. bounded. ∎
@@
For the reverse implication, for any $\xi\in L_{++}^{0}(\mathcal{F})$, let $T_{\xi}$ be the isomorphism of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ to itself, defined by $T_{\xi}(x)=\xi x$ for any $x\in L^{0}(\mathcal{F},\mathbb{R}^{d})$. Then it is also clear that $T_{\xi}^{-1}=T_{\xi^{-1}}$. Now, let $G$ and $f$ be the same as assumed in Proposition 2.3. Then there exists some $\xi\in L_{++}^{0}(\mathcal{F})$ such that $\xi^{-1}G\subseteq L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$, where $\bar{B}_{1}(\mathbb{R}^{d})=\{y\in\mathbb{R}^{d}:|y|\leq 1\}$ and $|\cdot|$ stands for the Euclidean norm on $\mathbb{R}^{d}$. For brevity, we denote $L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ by $V$, then $V$ is an a.s. bounded, $\mathcal{T}_{\varepsilon,\lambda}$-closed and $L^{0}$-convex subset of $L^{0}(\mathcal{F},\mathbb{R}^{d})$. Further, let $P_{V,\xi^{-1}G}$ be the restriction of $P_{\xi^{-1}G}$ to $V$, where $P_{\xi^{-1}G}$ is the random projection of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ onto $\xi^{-1}G$. Then $h:=T_{\xi^{-1}}\circ f\circ T_{\xi}\circ P_{V,\xi^{-1}G}$ is a $\sigma$-stable and $\mathcal{T}_{\varepsilon,\lambda}$-continuous mapping from $V$ to $V$, and hence $h$ is also a local and continuous-in-probability operator. Now, define a set-valued function $U:\Omega\rightarrow 2^{\mathbb{R}^{d}}$ by $U(\omega)=\bar{B}_{1}(\mathbb{R}^{d})$ for each $\omega\in\Omega$. Then $U$ not only satisfies the hypothesis of Proposition 2.4 but also is weakly measurable in the sense of [55], and in particular $G_{U}$ is exactly $V$. Thus, there exists $v\in V$ such that $h(v)=v$ by Proposition 2.4. Then $\xi v=f(\xi P_{V,\xi^{-1}G}(v))\in G$. Again, let $g\in G$ be such that $g=\xi v$. Then $v=\xi^{-1}g\in\xi^{-1}G$, and $P_{V,\xi^{-1}G}(v)=\xi^{-1}g$. Hence $f(\xi P_{V,\xi^{-1}G}(v))=f(g)=\xi v=g$, namely, $g$ is a fixed point of $f$ in $G$. ∎
@@
Let $V=L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ and $h:V\rightarrow V$ be the same as obtained in the process of the proof of that Proposition 2.4 implies Proposition 2.3. If $(\Omega,\mathcal{F},P)$ is a complete probability space, then by Lemma 5.2 there exists a Carathéodory function $f:\Omega\times\bar{B}_{1}(\mathbb{R}^{d})\rightarrow\bar{B}_{1}(\mathbb{R}^{d})$ such that $h(v)=h_{f}(v)$ for any $v\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$. This observation leads us to the following interesting (or even somewhat surprising) result.
@@
    Every Carathéodory function function $f:(\Omega,\mathcal{F},P)\times\bar{B}_{1}(\mathbb{R}^{d})\rightarrow\bar{B}_{1}(\mathbb{R}^{d})$ admits a random fixed point, namely, there exists a random vector $v:(\Omega,\mathcal{F},P)\rightarrow\bar{B}_{1}(\mathbb{R}^{d})$ such that $f(\omega,v(\omega))=v(\omega)$ a.s..
@@
$(1)\Rightarrow(2)$. Let $h=h_{f}$. Then there exists some $x\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ such that $h_{f}(x)=x$, and hence taking $v$ to be an arbitrarily chosen representative of $x$ yields $f(\omega,v(\omega))=v(\omega)$ a.s., namely, $v$ satisfies (2).
$(2)\Rightarrow(1)$. Let $\hat{\mathcal{F}}$ be the completion of $\mathcal{F}$ with respect to $P$ and further notice the fact that $L^{0}(\hat{\mathcal{F}},\bar{B}_{1}(\mathbb{R}^{d}))=L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ although $\hat{\mathcal{F}}$ and $\mathcal{F}$ may differ. Then by Lemma 5.2 there exists some Carathéodory function $f:(\Omega,\hat{\mathcal{F}},P)\times\bar{B}_{1}(\mathbb{R}^{d})\rightarrow\bar{B}_{1}(\mathbb{R}^{d})$ such that $h=h_{f}$. By (1) there exists some $^\hat{\mathcal{F}}$-measurable $v:\Omega\rightarrow\bar{B}_{1}(\mathbb{R}^{d})$ such that $f(\omega,v(\omega))=v(\omega)$ a.s.. Let $x$ be the equivalence class of $v$. Then $h(x)=h_{f}(x)=x$, since $L^{0}(\hat{\mathcal{F}},\bar{B}_{1}(\mathbb{R}^{d}))=L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$, $x$ is also regarded as a fixed point of $h$ in $L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$. ∎
@@
Let $G$ be a $\mathcal{T}_{\varepsilon,\lambda}$-closed $L^{0}$-convex subset of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ and $T:G\rightarrow L^{0}(\mathcal{F},\mathbb{R}^{d})$ be a $\sigma$-stable $\mathcal{T}_{\varepsilon,\lambda}$-continuous mapping. Then $T$ is a.s. sequentially continuous.
@@
Since elements of both $G$ and $L^{0}(\mathcal{F},\mathbb{R}^{d})$ are equivalence classes, we can, without loss of generality, assume that $(\Omega,\mathcal{F},P)$ is complete. Further, let $P_{G}$ be the random projection of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ onto $G$. Then $T\circ P_{G}:L^{0}(\mathcal{F},\mathbb{R}^{d})\rightarrow L^{0}(\mathcal{F},\mathbb{R}^{d})$ is still $\sigma$-stable and $\mathcal{T}_{\varepsilon,\lambda}$-continuous, and hence by Lemma 5.2 there exists some Carathéodory function $f:(\Omega,\mathcal{F},P)\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ such that $T\circ P_{G}=h_{f}$, in particular $T(g)=h_{f}(g)$ for any $g\in G$, which thus implies that $T$ is a.s. sequentially continuous. ∎
@@
6 The equivalence between the random Brouwer fixed point theorem and random Borsuk theorem
Proposition 6.1 below shows that Theorem 2.1 is equivalent to a special case of it, namely, random Brouwer fixed point theorem for random closed balls, which will provide much convenience for our research into the random version of the classical Borsuk theorem.
@@ 6 The equivalence between the random Brouwer fixed point theorem and random Borsuk theorem
    Every $\mathcal{T}_{c}$-continuous $\sigma$-stable mapping from a $\mathcal{T}_{c}$-closed, $\sigma$-stable, a.s. bounded and $L^{0}$-convex subset $G$ of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ to $G$ has a fixed point.
@@
    Every $\mathcal{T}_{c}$-continuous $\sigma$-stable mapping from a random closed ball $\bar{B}(\theta,r)$ of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ to $\bar{B}(\theta,r)$ has a fixed point, where $r\in L^{0}_{++}(\mathcal{F})$ and $\bar{B}(\theta,r)=\{x\in L^{0}(\mathcal{F},\mathbb{R}^{d}):\|x\|\leq r\}$.
@@
    Every $\mathcal{T}_{c}$-continuous $\sigma$-stable mapping from the random closed unit ball $\bar{B}(\theta,1)$ of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ to $\bar{B}(\theta,1)$ has a fixed point.
@@
$(2)\Rightarrow(1)$. Let $f:G\rightarrow G$ be a $\mathcal{T}_{c}$-continuous $\sigma$-stable mapping. Since $G$ is a.s. bounded, we can choose a sufficiently large $r\in L^{0}_{++}(\mathcal{F})$ such that $G\subseteq\bar{B}(\theta,r)$. Further, since $G$ is $\sigma$-stable and $\mathcal{T}_{c}$-closed, it follows immediately from [24] that $G$ is also $\mathcal{T}_{\varepsilon,\lambda}$-closed. Then by Lemma 4.12 the random projection $P_{G}$ of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ onto $G$ is well defined, $\sigma$-stable and $\mathcal{T}_{c}$-continuous. Now, let $P_{\bar{B}(\theta,r),G}$ be the restriction of $P_{G}$ to $\bar{B}(\theta,r)$. Then it is also $\sigma$-stable and $\mathcal{T}_{c}$-continuous since $\bar{B}(\theta,r)$ is $\sigma$-stable. Thus, it is clear that $f\circ P_{\bar{B}(\theta,r),G}$ is a $\mathcal{T}_{c}$-continuous and $\sigma$-stable mapping from $\bar{B}(\theta,r)$ to $\bar{B}(\theta,r)$. By (2) it follows that $f\circ P_{\bar{B}(\theta,r),G}$ has a fixed point $x$. Eventually, $f(x)=x$ since $x\in G$.
$(2)\Leftrightarrow(3)$ is clear since $T_{r}:\bar{B}(\theta,1)\rightarrow\bar{B}(\theta,r)$, defined by $T_{r}(x)=rx$ for each $x\in\bar{B}(\theta,1)$, is a $\sigma$-stable $\mathcal{T}_{c}$-homeomorphism. ∎
@@
Let $E$ and $F$ be the same as in Definition 6.2. If $h:F\rightarrow E$ is a $\sigma$-stable and bijective mapping, then it is easy to check that $h^{-1}:E\rightarrow F$ is also $\sigma$-stable. It is also clear that a $\sigma$-stable $\mathcal{T}_{c}$-homeomorphism (resp., $\mathcal{T}_{\varepsilon,\lambda}$-homeomorphism) of $F$ onto $E$ must be a $\mathcal{T}_{c}$-$r$-mapping (resp., a $\mathcal{T}_{\varepsilon,\lambda}$-$r$-mapping) of $F$ onto $E$. It follows immediately from [27, Lemma 4.3] that a $\mathcal{T}_{\varepsilon,\lambda}$-$r$-mapping (resp., a $\mathcal{T}_{\varepsilon,\lambda}$-retraction) of $F$ onto $E$ must be a $\mathcal{T}_{c}$-$r$-mapping (resp., a $\mathcal{T}_{c}$-retraction) of $F$ onto $E$.
Proposition 6.3 below is a random version of [1, Theorem 6.9], which is essentially a reformulation of Theorem 2.1 in a more general form.
@@
Let $g:E\rightarrow F$ and $h:F\rightarrow E$ be the same as in Definition 6.2. Then $g\circ f\circ h:F\rightarrow F$ has a fixed point $z$ by Theorem 2.1. Let $x=h(z)$. Then
@@
Theorem 6.4 below can be aptly called a random Borsuk theorem, which is a random version of the classical Borsuk theorem [2] (see also [1, Theorem 6.12]). Here, we would like to remind the reader that $L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ in the remainder of this section is the very same thing as the random closed unit ball $\bar{B}(\theta,1):=\{x\in L^{0}(\mathcal{F},\mathbb{R}^{d}):\|x\|\leq 1\}$.
@@
$L^{0}(\mathcal{F},\partial\bar{B}_{1}(\mathbb{R}^{d}))$ is not a $\mathcal{T}_{c}$-$r$-image of $L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$, where, as usual, $\bar{B}_{1}(\mathbb{R}^{d})=\{y\in\mathbb{R}^{d}:|y|\leq 1\}$, $\partial\bar{B}_{1}(\mathbb{R}^{d})=\{y\in\mathbb{R}^{d}:|y|=1\}$ and $|\cdot|$ denotes the Euclidean norm on $\mathbb{R}^{d}$. In particular, $L^{0}(\mathcal{F},\partial\bar{B}_{1}(\mathbb{R}^{d}))$ is not a $\mathcal{T}_{c}$-retraction of $L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$.